# Uncovering the first-infall history of the LMC through its dynamical impact in the Milky Way halo

### v1

@@ 4.5 Compare observation with simulations
We now proceed to compare the observed latitudinal velocities ($v_{b}$) from the RR Lyrae sample with the outcomes derived from our grid of MW-LMC simulations. Our previous analysis of the LMC-induced perturbations was carried out in the Galactocentric coordinate system within the context of N-body simulations, where the observer is placed at the center of the MW. This perspective provides us with a clearer understanding of the LMC’s impact on Galactic halo and facilitates the identification of asymmetries in the density and velocity fields. However, given that our RR Lyrae sample is limited to 5-dimensional phase-space coordinates (lacking the radial velocity data), we are unable to calculate their velocities in the Galactocentric frame. Instead, we focus on their mean latitudinal velocities within the Galactic heliocentric coordinate system, where the observer is located at (8.30, 0.00, 0.02) kpc in the Galactocentric cartesian frame (Bennett & Bovy, 2019) and is at rest with respect to the sun, which orbits around the Galactic center at a velocity of (11.10, 244.24, 7.25) km/s (Schönrich et al., 2010; McMillan, 2017).
@@
Figure 9 demonstrates the mean latitudinal velocity $\langle v_{b}\rangle$ in the heliocentric frame. The $x$-axis represents the virial mass of the MW halo, which spans from $0.5\times 10^{12}\mathrm{M}_{\odot}$ to $1.8\times 10^{12}\mathrm{M}_{\odot}$, while the $y$-axis shows the average latitudinal velocity for both the RR Lyrae sample and the simulated star catalog. We estimate the uncertainty of RR Lyrae stars by Monte-Carlo sampling the errors in their observables (i.e. proper motions measurements from Gaia and distances estimates from Li et al. (2023)) 10,000 times. For simplicity, we do not introduce measurement covariances as the distance and proper motion measurements in our study are done independently, and we deem the impact of the proper motion covariances on our analysis to be small. While for the statistical values in simulations, as elaborated in the Section 4.4, we first introduce realistic measurement uncertainties in stars’ distances and velocities to the present-day simulation snapshots, and then derive the mean velocities and their uncertainties using the bootstrap technique by repeatedly drawing samples of particles with replacement from the simulation, with each bootstrap sample having the same particle numbers and the spatial distribution as the 1,165 RR Lyrae sample. We find that the uncertainties in our simulated catalogs are larger than those in the observational data. This discrepancy arises because our estimates account for both the intrinsic noise introduced into the distances and velocities of particles, and additional statistical sampling noise generated when we randomly select particles to create bootstrap samples.
@@
The blue shaded region in the figure represents MW-LMC systems with a LMC mass of $1.5\times 10^{11}\mathrm{M}_{\odot}$ before it enters into the MW, a spherical Galactic halo with an isotropic velocity profile, but varying MW halo masses. In the heliocentric frame, the mean latitudinal velocities $\langle v_{b}\rangle$ for these models are about $25\pm 5$ km/s in the first-infall scenario (represented by the solid line), where the MW halo mass ranges from $0.5\times 10^{12}\mathrm{M}_{\odot}$ to $1.0\times 10^{12}\mathrm{M}_{\odot}$, and decrease to $18\pm 5$ km/s in the second-passage scenario (represented by the dashed line), with the halo mass ranging from $1.1\times 10^{12}\mathrm{M}_{\odot}$ to $1.5\times 10^{12}\mathrm{M}_{\odot}$. The red shaded region represents models with the same MW potentials and velocity profile as the blue shaded region but with a heavier LMC mass at infall ($M_{\text{LMC}}=2.1\times 10^{11}\mathrm{M}_{\odot}$). In these models, the mean latitudinal velocities increase by about 3 km/s to $28\pm 5$ km/s in the first-infall scenario and by 1 km/s to $19\pm 5$ km/s in the second-passage scenario.
@@
In addition to the models featuring a MW halo with isotropic velocity profile ($\beta(r)$ = 0), the green shaded region maintains the same MW and LMC potentials as the blue shaded region but assumes a MW halo with radially varying anisotropic velocity profile (Hansen & Moore, 2006): $\beta(r)=-0.15-0.2\alpha(r)$, where $\beta(r)$ is defined as $\beta(r)=1-\frac{\sigma_{t}(r)^{2}}{2\sigma_{r}(r)^{2}}$, with $\sigma_{r}$ and $\sigma_{t}$ being the radial and tangential velocity dispersion of halo stars, respectively, and $\alpha(r)=\frac{\mathrm{d}\ln\rho(r)}{d\ln r}$. This choice is motivated by previous findings indicating that a radially varying anisotropic velocity profile can alter the morphology of the LMC-induced dynamical effects observed in the sky (Garavito-Camargo et al., 2019; Vasiliev, 2023a). In this case, $\langle v_{b}\rangle$ is about $22\pm 5$ km/s in the first-infall scenario, lower by about 3 km/s compared to the blue shaded region, and decreases to $17\pm 5$ km/s in the second-passage scenario.
@@
Moreover, the purple shaded region represents scenarios where the spherical MW halo transitions to an oblate shape ($q$ = 0.7), with its minor axis perpendicular to the Galactic disk. This change increases the transition MW halo mass between the first-infall and second-passage scenarios of the LMC’s orbital history to $1.5\times 10^{12}\mathrm{M}_{\odot}$ and reduces the magnitudes of $\langle v_{b}\rangle$ in both scenarios to lower values compared to previous cases. Specifically, $\langle v_{b}\rangle$ is about $14\pm 5$ km/s in the first-infall scenario, where the MW halo mass ranges from $1.0\times 10^{12}\mathrm{M}_{\odot}$ to $1.4\times 10^{12}\mathrm{M}_{\odot}$, and about $8\pm 5$ km/s in the second-passage scenario, with the halo mass ranging from $1.5\times 10^{12}\mathrm{M}_{\odot}$ to $1.8\times 10^{12}\mathrm{M}_{\odot}$.
@@
As a result, our analysis of the kinematic properties of the RR Lyrae sample (yellow shaded region), which has a mean latitudinal velocity $\langle v_{b}\rangle=30.8\pm 4.0$ km/s supports the first-infall scenario as the likely orbital history for the LMC. In particular, for the MW-LMC system where the LMC has a mass of $2.1\times 10^{11}\mathrm{M}_{\odot}$ at infall, the LMC-induced perturbations in the first-infall branch (with $M_{\text{MW}}\leq 1.0\times 10^{12}\mathrm{M}_{\odot}$) have mean latitudinal velocities of approximately $28\pm 5$ km/s. This is more consistent with observational data than scenarios involving a less massive LMC ($M_{\text{LMC}}=1.5\times 10^{11}\mathrm{M}_{\odot}$), where the observed mean latitudinal velocity is $\langle v_{b}\rangle\leq 25\pm 5$ km/s, regardless of the MW halo mass and velocity profile. Therefore, future studies on the MW-LMC interaction should consider LMC models with a virial mass close to $2.1\times 10^{11}\mathrm{M}_{\odot}$ at the time of infall. Furthermore, in simulations with an oblate MW halo ($q$ = 0.7), both the first-infall and second-passage branches fall below the average latitudinal velocity in observation. This suggests that, for the simplified, non-tilted, spheroidal MW model aligned with the Galactic disk, the flattening parameter $q$ should be larger than 0.7. For example, if the MW halo has a flattening parameter of 0.8, it’s plausible that the mean latitudinal velocities in the first-infall scenario (with $M_{\text{MW}}\leq 1.3\times 10^{12}\mathrm{M}_{\odot}$, for further discussion on the degeneracy between the transition MW mass and the flattening parameter, refer to section 5.6) can align with observational data when the LMC has a large virial mass at infall.
@@ 5.1 Comparison to Previous Work
As for the shape of the MW halo. The majority of previous works report an oblate spheroid halo that is aligned with the Galactic disk and the flattening parameter $q$ is constrained to values close to 1.0 in some recent works. Hernitschek et al. (2018) use 11,000 RRab stars (a subtype of RR Lyrae star) within $20\mathrm{kpc}\leq R_{\mathrm{gc}}\leq 131\mathrm{kpc}$ from Pan-STARRS1 to measure the density profile of the Galactic halo. The distribution of the sources is then well fit by a ellipsoidal stellar density model with a constant flattening parameter $q$ and $q=0.918_{-0.014}^{+0.016}$. In addition, Hattori et al. (2021) model the stellar halo distribution function with the astrometric data for halo RR Lyrae stars within 30 kpc from Gaia DR2 and find a posterior distribution of $q>0.963$. These measurements are consistent with our results obtained from the kinematic perturbations induced by the LMC in the MW halo, which favors $q>0.7$.
However, the shape of the real MW halo can be more complex than the simple model, which is parametrized by a single virial mass and $q$. For example, the flattening parameter $q$ can be different in the inner and outer halo. Hernitschek et al. (2018) also fit their star sample with a ellipsoidal stellar density model with a radius-dependent flattening parameter and find the stellar halo is more flattened ($q\sim 0.8$) at inner halo (r $\sim 25$ kpc) while more spherical ($q\sim$ 1.0) at outer halo. Furthermore, MW halo may tilt relative to the disk plane and present a triaxial shape. Han et al. (2022a) use 5559 halo stars belonging to GSE within the radial range 6-60 kpc to infer the density profile of the accreted stellar halo. They find the shape of the accreted halo can be fit by a tilted, triaxial ellipsoid with the principal axes having a ratio of 10:8:7 and the major axis pointing 25 degree above the Galactic plane towards the position of the sun. Extending the analysis beyond our simple parametrized models will be crucial for investigating these subtle properties of the MW halo.
@@ 5.5 Possible stellar halo sloshing from GSE
Figure 12: The ongoing sloshing of GSE on radial and latitudinal velocities within two radii intervals (30-60 kpc and 60-90 kpc) in the Galactocentric frame. Our analysis is conducted through a similar N-body simulation of the MW-GSE merger as described in Naidu et al. (2021). We find that the remaining sloshing in both radial and latitudinal velocities shows average values of -1.42 km/s and 0.09 km/s for $30\mathrm{kpc}<R_{\text{gal}}<60\mathrm{kpc}$, and -3.38 km/s and -0.36 km/s for $60\mathrm{kpc}<R_{\text{gal}}<90\mathrm{kpc}$, respectively. The magnitudes of these perturbations are negligible compared to the predicted influences of the LMC if it follows a first-infall orbital scenario.
@@
The results of the present-day simulation snapshot are shown in Figure 12. Similar to Figure 4-6, we illustrate the influence of GSE on radial and latitudinal velocities within two radii intervals (30-60 kpc and 60-90 kpc) in Galactocentric frame. We find that the remaining sloshing in both radial and latitudinal velocities shows average values of -1.42 km/s and 0.09 km/s for $30\mathrm{kpc}<R_{\text{gal}}<60\mathrm{kpc}$, and -3.38 km/s and -0.36 km/s for $60\mathrm{kpc}<R_{\text{gal}}<90\mathrm{kpc}$, respectively. Compared to the three representative cases demonstrated in Figures 4-6, when the LMC follows a first-infall orbital scenario, the residual sloshing from the GSE, especially in the latitudinal velocity, is negligible compared to the predicted influences of the LMC. However, if the LMC is in the second-passage scenario, these residual effects from the GSE may be comparable on the LMC-induced perturbations in radial velocities. Given that our study supports the first-infall scenario for the LMC, the potential impact of stellar sloshing from the GSE on the LMC-induced kinematic perturbations is minimal.
@@ 5.7 Caveats and future directions
In this study, we also primarily compare the mean latitudinal velocities of RR Lyrae stars in the Galactic frame with those from our simulated star catalogs, owing to the availability of 5-dimensional phase-space information for distant halo stars. However, our simulations suggest that (see Figure 8) the radial velocities of halo stars in the southern hemisphere of the Milky Way can serve as an additional effective indicator of the LMC’s orbital history. Unlike tangential velocities, radial velocities of halo stars can be measured with a higher degree of precision in ongoing and forthcoming spectroscopic surveys such as WEAVE (Jin et al., 2023), 4MOST (de Jong et al., 2016), DESI (Levi et al., 2019) and PFS (Greene et al., 2022). The availability of complete 6-dimensional phase-space information for these halo stars also enables us to investigate the perturbations in the Galactocentric frame, providing a robust view of the intrinsic density and kinematic properties of the MW halo.

### v2

@@ 4.5 Compare observation with simulations
We now proceed to compare the observed latitudinal velocities ($v_{b}$) from the RR Lyrae sample with the outcomes derived from our grid of MW-LMC simulations. Our previous analysis of the LMC-induced perturbations was carried out in the Galactocentric coordinate system within the context of N-body simulations, where the observer is placed at the center of the MW. This perspective provides us with a clearer understanding of the LMC’s impact on Galactic halo and facilitates the identification of asymmetries in the density and velocity fields. Given that our RR Lyrae sample is limited to 5-dimensional phase-space coordinates (lacking the radial velocity data), we use reflex-correct function from gala package to correct for the solar reflex motion in the stars’ proper motions and calculate their velocities in the Galactocentric frame. For this correction, we set the position and velocity of the Sun in the Galactocentric cartesian frame to (8.30, 0.00, 0.02) kpc (Bennett & Bovy, 2019) and (11.10, 244.24, 7.25) km/s (Schönrich et al., 2010; McMillan, 2017) respectively.
@@
Figure 9 demonstrates the mean latitudinal velocity $\langle v_{b}\rangle$ in the Galactocentric frame. The $x$-axis represents the virial mass of the MW halo, which spans from $0.5\times 10^{12}\mathrm{M}_{\odot}$ to $1.8\times 10^{12}\mathrm{M}_{\odot}$, while the $y$-axis shows the average latitudinal velocity for both the RR Lyrae sample and the simulated star catalog. We estimate the uncertainty of RR Lyrae stars by Monte-Carlo sampling the errors in their observables (i.e. proper motions measurements from Gaia and distances estimates from Li et al. (2023)) 10,000 times. For simplicity, we do not introduce measurement covariances as the distance and proper motion measurements in our study are done independently, and we deem the impact of the proper motion covariances on our analysis to be small. While for the statistical values in simulations, as elaborated in the Section 4.4, we first introduce realistic measurement uncertainties in stars’ distances and velocities to the present-day simulation snapshots, and then derive the mean velocities and their uncertainties using the bootstrap technique by repeatedly drawing samples of particles with replacement from the simulation, with each bootstrap sample having the same particle numbers and the spatial distribution as the 1,126 RR Lyrae sample. We find that the uncertainties in our simulated catalogs are larger than those in the observational data. This discrepancy arises because our estimates account for both the intrinsic noise introduced into the distances and velocities of particles, and additional statistical sampling noise generated when we randomly select particles to create bootstrap samples.
@@
The blue shaded region in the figure represents MW-LMC systems with a LMC mass of $1.5\times 10^{11}\mathrm{M}_{\odot}$ before it enters into the MW, a spherical Galactic halo with an isotropic velocity profile, but varying MW halo masses. The mean latitudinal velocities $\langle v_{b}\rangle$ for these models are about $13\pm 4$ km/s in the first-infall scenario (represented by the solid line), where the MW halo mass ranges from $0.5\times 10^{12}\mathrm{M}_{\odot}$ to $1.0\times 10^{12}\mathrm{M}_{\odot}$, and decrease to below $7\pm 4$ km/s in the second-passage scenario (represented by the dashed line), with the halo mass ranging from $1.1\times 10^{12}\mathrm{M}_{\odot}$ to $1.5\times 10^{12}\mathrm{M}_{\odot}$. The red shaded region represents models with the same MW potentials and velocity profile as the blue shaded region but with a heavier LMC mass at infall ($M_{\text{LMC}}=2.1\times 10^{11}\mathrm{M}_{\odot}$). In these models, the mean latitudinal velocities increase by about 3 km/s to $16\pm 4$ km/s in the first-infall scenario and by 1 km/s to below $8\pm 4$ km/s in the second-passage scenario.
@@
In addition to the models featuring a MW halo with isotropic velocity profile ($\beta(r)$ = 0), the green shaded region maintains the same MW and LMC potentials as the blue shaded region but assumes a MW halo with radially varying anisotropic velocity profile (Hansen & Moore, 2006): $\beta(r)=-0.15-0.2\alpha(r)$, where $\beta(r)$ is defined as $\beta(r)=1-\frac{\sigma_{t}(r)^{2}}{2\sigma_{r}(r)^{2}}$, with $\sigma_{r}$ and $\sigma_{t}$ being the radial and tangential velocity dispersion of halo stars, respectively, and $\alpha(r)=\frac{\mathrm{d}\ln\rho(r)}{d\ln r}$. This choice is motivated by previous findings indicating that a radially varying anisotropic velocity profile can alter the morphology of the LMC-induced dynamical effects observed in the sky (Garavito-Camargo et al., 2019; Vasiliev, 2023a). In this case, $\langle v_{b}\rangle$ is about $12\pm 4$ km/s in the first-infall scenario, lower by about 1 km/s compared to the blue shaded region, and decreases to below $7\pm 4$ km/s in the second-passage scenario.
@@
Moreover, the purple shaded region represents scenarios where the spherical MW halo transitions to an oblate shape ($q$ = 0.7), with its minor axis perpendicular to the Galactic disk. This change increases the transition MW halo mass between the first-infall and second-passage scenarios of the LMC’s orbital history to $1.5\times 10^{12}\mathrm{M}_{\odot}$ and reduces the magnitudes of $\langle v_{b}\rangle$ in both scenarios to lower values compared to previous cases. Specifically, $\langle v_{b}\rangle$ is about $12\pm 4$ km/s in the first-infall scenario, where the MW halo mass ranges from $1.0\times 10^{12}\mathrm{M}_{\odot}$ to $1.4\times 10^{12}\mathrm{M}_{\odot}$, and about $4\pm 4$ km/s in the second-passage scenario, with the halo mass ranging from $1.5\times 10^{12}\mathrm{M}_{\odot}$ to $1.8\times 10^{12}\mathrm{M}_{\odot}$.
@@
As a result, our analysis of the kinematic properties of the RR Lyrae sample (yellow shaded region), which has a mean latitudinal velocity $\langle v_{b}\rangle=18.1\pm 4.1$ km/s supports the first-infall scenario as the likely orbital history for the LMC. This value is consistent with a recent measurement from Chandra et al. (2024) using red giant stars. In particular, for the MW-LMC system where the LMC has a mass of $2.1\times 10^{11}\mathrm{M}_{\odot}$ at infall, the LMC-induced perturbations in the first-infall branch (with $M_{\text{MW}}\leq 1.0\times 10^{12}\mathrm{M}_{\odot}$) have mean latitudinal velocities of approximately $16\pm 4$ km/s. This aligns more closely with observational data than scenarios involving a less massive LMC ($M_{\text{LMC}}=1.5\times 10^{11}\mathrm{M}_{\odot}$), where the observed mean latitudinal velocity is $\langle v_{b}\rangle\approx 13\pm 4$ km/s, regardless of the MW halo mass and velocity profile. Therefore, future studies on the MW-LMC interaction should consider LMC models with a virial mass close to $2.1\times 10^{11}\mathrm{M}_{\odot}$ at the time of infall. Furthermore, in simulations with an oblate MW halo ($q$ = 0.7), both the first-infall and second-passage branches fall below the average latitudinal velocity in observation. This suggests that, for the simplified, non-tilted, spheroidal MW model aligned with the Galactic disk, the flattening parameter $q$ should be larger than 0.7. For example, if the MW halo has a flattening parameter of 0.8, it’s plausible that the mean latitudinal velocities in the first-infall scenario (with $M_{\text{MW}}\leq 1.3\times 10^{12}\mathrm{M}_{\odot}$, for further discussion on the degeneracy between the transition MW mass and the flattening parameter, refer to section 5.6) can align with observational data when the LMC has a large virial mass at infall.
@@ 5.1 Comparison to Previous Work
As for the shape of the MW halo. The majority of previous works report an oblate spheroid halo that is aligned with the Galactic disk and the flattening parameter $q$ is constrained to values close to 1.0 in some recent works. Küpper et al. (2015) compared synthetic models of the Pal 5 stream’s tidal tails with locally overdense regions of selected stars from SDSS DR9, finding that the halo of the MW has a flattening parameter of $q=0.95_{-0.12}^{+0.16}$. Hernitschek et al. (2018) use 11,000 RRab stars (a subtype of RR Lyrae star) within $20\mathrm{kpc}\leq R_{\mathrm{gc}}\leq 131\mathrm{kpc}$ from Pan-STARRS1 to measure the density profile of the Galactic halo. The distribution of the sources is then well fit by a ellipsoidal stellar density model with a constant flattening parameter $q$ and $q=0.918_{-0.014}^{+0.016}$. In addition, Hattori et al. (2021) model the stellar halo distribution function with the astrometric data for halo RR Lyrae stars within 30 kpc from Gaia DR2 and find a posterior distribution of $q>0.963$. These measurements are consistent with our results obtained from the kinematic perturbations induced by the LMC in the MW halo, which favors $q>0.7$.
However, the shape of the real MW halo can be more complex than our simple model, which is characterized by a single flattening parameter $q$. For example, $q$ can be different in the inner and outer halo. Hernitschek et al. (2018) fit their star sample with a ellipsoidal stellar density model with a radius-dependent flattening parameter and find the stellar halo is more flattened ($q\sim 0.8$) at inner halo (r $\sim 25$ kpc) while more spherical ($q\sim$ 1.0) at outer halo. Furthermore, MW halo may tilt relative to the stellar disk plane and present a triaxial shape. This has been demonstrated by fitting the phase-space structures of tidal streams spanning large arcs in the sky. Using the Sagittarius stream, Law & Majewski (2010) and Vasiliev et al. (2021) introduce a non-axisymmetric potential/density model, which features a nearly oblate ($q\sim 0.7$) outer halo with its minor axis lying in the disc plane and pointing approximately in the direction of the LMC’s orbital pole. Meanwhile, considering a different halo model with an axisymmetric shape and a radially constant flattening in the potential, Erkal et al. (2019) also find that the direction of flattening is roughly aligned with the orbital plane of the LMC through the orbit-fitting of the Orphan stream. Beyond evidence from large-scale stellar streams, Han et al. (2022a) use 5559 halo stars belonging to GSE within the radial range 6-60 kpc to infer the density profile of the accreted stellar halo. They find the shape of the accreted halo can be fit by a tilted, triaxial ellipsoid with the principal axes having a ratio of 10:8:7 and the major axis pointing 25 degree above the Galactic plane towards the position of the sun. Extending the analysis beyond our simple parametrized models in future studies will be crucial for investigating these subtle properties of the MW halo.
@@ 5.5 Possible stellar halo sloshing from GSE
Figure 12: The ongoing sloshing of GSE on radial and latitudinal velocities within two radii intervals (30-60 kpc and 60-90 kpc) in the Galactocentric frame. Our analysis is conducted through a similar N-body simulation of the MW-GSE merger as described in Naidu et al. (2021). We find that the remaining sloshing in both radial and latitudinal velocities shows average values of -0.86 km/s and -0.23 km/s for $30\mathrm{kpc}<R_{\text{gal}}<60\mathrm{kpc}$, and -2.14 km/s and 0.34 km/s for $60\mathrm{kpc}<R_{\text{gal}}<90\mathrm{kpc}$, respectively. The magnitudes of these perturbations are negligible compared to the predicted influences of the LMC if it follows a first-infall orbital scenario.
@@
The results of the present-day simulation snapshot are shown in Figure 12. Similar to Figure 4-6, we illustrate the influence of GSE on radial and latitudinal velocities within two radii intervals (30-60 kpc and 60-90 kpc) in Galactocentric frame. We find that the remaining sloshing in both radial and latitudinal velocities shows average values of -0.86 km/s and -0.23 km/s for $30\mathrm{kpc}<R_{\text{gal}}<60\mathrm{kpc}$, and -2.14 km/s and 0.34 km/s for $60\mathrm{kpc}<R_{\text{gal}}<90\mathrm{kpc}$, respectively. Compared to the three representative cases demonstrated in Figures 4-6, when the LMC follows a first-infall orbital scenario, the residual sloshing from the GSE, especially in the latitudinal velocity, is negligible compared to the predicted influences of the LMC. However, if the LMC is in the second-passage scenario, these residual effects from the GSE may be comparable on the LMC-induced perturbations in radial velocities. Given that our study supports the first-infall scenario for the LMC, the potential impact of stellar sloshing from the GSE on the LMC-induced kinematic perturbations is minimal.
@@ 5.7 Caveats and future directions
In this study, we also primarily compare the mean latitudinal velocities of RR Lyrae stars in the Galactocentric frame with those from our simulated star catalogs, owing to the availability of 5-dimensional phase-space information for distant halo stars. However, our simulations suggest that (see Figure 8) the radial velocities of halo stars in the southern hemisphere of the Milky Way can serve as an additional effective indicator of the LMC’s orbital history. Unlike tangential velocities, radial velocities of halo stars can be measured with a higher degree of precision in ongoing and forthcoming spectroscopic surveys such as WEAVE (Jin et al., 2023), 4MOST (de Jong et al., 2016), DESI (Levi et al., 2019) and PFS (Greene et al., 2022). The availability of complete 6-dimensional phase-space information for these halo stars also enables us to investigate the perturbations in the Galactocentric frame, providing a robust view of the intrinsic density and kinematic properties of the MW halo.